Find the sum of the series 1+3.5+6+8.5+......+101
Find the sum of the first 23 term of the AP 4-3-10.....
An arithmetic series have 1st term as 4 and common difference as 1/2 find the first 20term
Find the sum of the first 100 term
Find the sum of arithmetic series term 1common difference 3 and last term 100
Find the sum of d first 20 term of arithmetic is identical to the sum of first 22 term. If the common difference is -2 find the first term
Osakpolormario@gmail.com

Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20056)

Find the sum of the series 1+3.5+6+8.5+......+101.
Sn = (n/2)(a1 + an)

a1 = 1

But we need first to determine n.
To do that we need d:

d = 3.5-1 = 6-3.5 = 8.5-6 = 2.5

an = a1 + (n - 1)∙d

101 = a1 + (n - 1)∙2.5

Solve that and get n = 41

Sn = (n/2)(a1 + an)

S41 = (41/2)(1 + a41)

S41 = (41/2)(1 + 101)

S41 = (41/2)(102)

S41 = 2091

------------------

Find the sum of the first 23 term of the AP 4-3-10.....
First find d:

d = -3 - 4 = -10 - (-3) = -7

Then use

Sn = (n/2)[2∙a1 + (n-1)∙d

with n = 23, a1 = 4, d = -7 

An arithmetic series have 1st term as 4 and common difference as 1/2 find the first 20 terms.
Write the first term as 4 and keep adding 1/2 over and over until
you have written 20 terms. To get you started,

the 2nd term is 4 + 1/2 = 8/2 + 1/2 = 9/2,
the 3rd term is 9/2 + 1/2 = 10/2 = 5,
etc., etc.,

Edwin

Answer by ikleyn(52788)

.
Hey,

welcome to the forum.


Please learn and accept some RULES of this forum regarding presenting your posts.

1.   One and only one problem/question per post.

2.   Use the comma and blank space as separators between the numbers in sequence.

3.   Be attentive and write your posts correctly.

        Check your post before submitting it.

        Remember:  it is your responsibility to present the info in the post by a correct way.

        It is not the tutors duties to guess what you were going to communicate in your post.


With it,  welcome to this forum !